Gundam 0083 Stardust Memory - Episode 12: Assault! Critical Intercept Point Page 1

Gundam 0083 Stardust Memory Episode 12: Assault! Critical Intercept Point

Translated By Tonghyun “Vajra” Kim (tkim@sdcc3.ucsd.edu) Edited by Noel Gamboa (egamboa@sdcc13.ucsd.edu) OPERATION X Version 1.0 11/16/92

Narration

The Albion docks at La Vienne Rose, to acquire the new Gundam Unit 3. But awaiting them was the Military Police Unit under the command of Lt. Commander Nakato. Ignoring the irritated Uraki and his people, Delaz ignites the propellants of the colony. The earth-bound colony… and fallen Lucette… Her death forces Nina to face her destiny. Now, Uraki enters the decisive battle to stop the colony from falling to the earth, with Gundam Unit 3.

Opening song

MEN OF DESTINY

Narration

Universal Century 0083 November 12. Operation Stardust, the purpose of which is to drop a colony on Earth, has entered its climax.

The Cima fleet, escorting the hijacked colony, wait to make contact with the Federation fleet. Without warning, Gundam Unit 3 Dendrobium Stamen approaches the convoy.

Helmsman

Enemy sighted! 10 o’clock portside!

Cima

In front of us?

Deatroaf

The Earth Orbital Fleet couldn’t have arrived yet.

Cima

I’ve heard nothing about it! I’m going to the comm room!

Hustler

This is Axis Fleet. I’m sorry. We’ll have to stand by and watch from the rear.

Delaz

That’s fine. You don’t need to worry about us. I leave the aftermath to you.

Hustler

I swear by the honor of Axis.

Delaz

It’s hard to believe we made it this far. With this fleet, I’m now doing the operation which required half of the Jion fleet in the past.

Unit 3 engages several Zakus and Doms.

Zaku pilot

He’s just big, that’s all.

Unit 3 makes short work of the Jion defense and continues its approach towards the colony. In General Kowen’s HQ…

Kowen

Status report.

Operator 1

Aye, sir. The colony’s at the halfway point between the moon and the earth. Gundam Unit 3 has made contact with it. And the Axis Fleet, which declared itself neutral, is retreating to the rear.

Kowen

Although they’re our enemies, this colony drop is an excellent one, making good use of the lessons of the One Year War.

Operator 1

But even Gundam Unit 3 can’t handle it alone.

Operator 2

Fourteen hours before it collides with the earth…

Kowen

Only nine hours before the game’s over. After that, the colony will pass the Critical Intercept Point.

Operator 2

Critical Intercept Point?

Operator 1

Here. If the colony goes beyond it, there’s no way we can stop something that heavy.

Kowen

We have to recover the colony and change its course before it reaches that point.

Operator 2

Will there be any remaining propellants?

Kowen

There must be! For the final course adjustment to Jaburo! Damn… What the hell is my Third Orbital Fleet doing?

Operator

They’re gathering at point 302 near the Critical Point to intercept the enemy.

Kowen

We won’t make it. What’s on the Admiral’s mind?

Back in space…

Kou

They’re retreating? Where is he?

Gato, aboard the mobile armor Neue Ziel, appears behind a Musai wreck.

Gato

I will not speak to him anymore.

Kou

It’s Gato!?

Gato

An I-Field!?

Tech note: An I-field deflects beams weapons by maintaining a field of similarly-charged particles.

Kou

Him too?

The Albion, trailing behind Gundam Unit 3, speeds towards the combat zone.

Sinapus

Distance?

Scott

107,000 to the colony.

Hallida

Captain! 8 hours and 22 minutes until the Critical Intercept Point. Please authorize mobile suits to scramble.

Sinapus

Not yet. Not until Unit 3 breaches the enemy defense.

Sinapus

Can Lt. Uraki’s body take the strain?

The Albion’s MS pilots await combat from inside their respective cockpits.

Adel

I never imagined that we’d get involved in this kind of incident.

Bait

Come to think of it, we’ve been haunted by the spectre of Jion.

Monsha

Ever since we met the Gundam…

Gato

With beams getting deflected like this…

Zaku pilot

Commander, with the unit armed mainly with beam weapons…

Gato

Don’t interfere!

Kou

He stopped moving?

Gato’s MA is rocked by Kou’s missiles.

Kou

Now!!!

Kou

The weapons - !

The Albion appears behind Unit 3.

Scott

Gundam, this is the Albion. We’ve caught up. Return to ship to resupply. Return to ship for now.

Kou

If I hang on, I can breach the first line of defense!

Sinapus

Lt. Uraki! You can’t recover the colony alone!

Kou

Roger.

Aboard Delaz’s flagship, the Gwaden.

Delaz

Well, the problem is when the enemy will appear on our front.

Staff

The Fed’s Earth Orbital Fleet is made up mostly of Salamis-class ships. There isn’t much to worry about.

Cima

But you couldn’t have imagined Gato getting thrown off-guard, could you?

Delaz

Unexpected things do happen. Gato is doing fine.

Cima

Unexpected things do happen…

The Federation fleet from Konpei, which had ran out of fuel earlier because of Delaz’s moon feint, prepares to resume pursuit of the hijacked colony.

Operator

This is the 73rd Fleet. Refueling completed. We’ll go ahead.

Konpei base commander

Admiral, we’ll resume pursuit. Even if we can’t catch up with the enemy, let us take care of the aftermath.

Konpei base commander

Huh? What did you just say?

Kou returns the Dendrobium Stamen to the Albion.

Mora

Begin replacing the weapon container! Take it out!

Mora

Nina…

A siren blares out!

Sinapus

A pursuit party? Pull back to the limit!

Kou

The enemy? We’ve been followed? Mora, I’m taking the Stamen out! Please clear the way.

Mora

No way! You’re in no condition to fight!

Monsha

They broke through us!

Kou

I’m going!

Nina

Men are really…

EYECATCH

Kowen pleads with the Admiral, who is accompanied by Jamitov Hyman. Cima’s grinning visage vanishes from a viewscreen.

Kowen

Admiral…

Jamitov

If we surround the enemy too tightly, they’ll fight to the death.

Kowen

I didn’t ask you! Admiral, what’s on your mind? No, what do you want to use this situation for? We should move the Earth Orbital Fleet to the front line…

Kowen finds a gun pointed at him.

Kowen

Jamitov, you bastard…

Kowen

We don’t have time. Please stop thinking about some strange scheme and order the fleet to attack…

Kowen

What the hell are you doing? Let me go! Admiral! We don’t have any time left!

Admiral

Why did Unit 3 fall into the Albion’s hands?

Out in space, the Neue Ziel is resupplied.

Gato

We can’t count on the likes of Cima Fleet. We’ll do it! With our own hands!

Gradol

But… your body won’t take the strain…

Gato

Gradol… And that Gundam… Only I can defeat him.

Gradol

Commander…

Gato

Never before I have looked at the earth with such excitement in my heart. I’ll defeat the Gundam before the enemy appears on our front.

Like Gato, Kou sits in his cockpit, wondering what the future holds.

Kou

Simmone… How many hours left?

Simmone

We just finished supplying.

Kou

No, I mean till the time limit.

Simmone

3 hours and 11 minutes.

Kou

Nina…

Sinapus

No time left already, and no idea when friendly forces will arrive… Nina Purpleton.

Nina

Yes?

Sinapus

I grant you permission to use the lifeboat. Please leave this ship immediately.

Nina

Capt. Sinapus, I want to see this till the very end with my own eyes, out of my own will. I’m sorry.

Sinapus

But the military won’t pay your salary.

Sinapus

Alright! Sail at second battle speed for five minutes then change to first battle speed!

Gato

Cima’s fleet has advanced too far. Understand? Be alert! Remember, we have nowhere to return to!

Gato is wondering why Cima’s ships are pushing too far ahead, but the answer becomes obvious by the end of this episode. Meanwhile, Cima has taken over Delaz’s ship, killing his bridge crew.

Deatroaf

Madame Cima, sensor readings of our front. It’s time.

Cima

Perfect timing.

Gato

Don’t go, Gato.

Scott

Captain, we picked up something portside! A light!

Sinapus

What? In Earth’s direction? Can you zoom in? Earth Orbital Fleet… No, what’s that? That array is…

Monsha

A wishbone formation!

A slew of Jion ships and mobile suits led by the Neue Ziel attack the lonely Albion!

Gato

The ship is not retreating. They must be determined.

Gato

It’s him?!

Kou

Gato!

Kou demolishes Gradol’s Musai battleship.

Gradol

Commander!

The Gwaden.

Cima

Distance to Critical Intercept Point?

Deatroaf

12000. 21 minutes to go.

Cima

After that point, even those annoying bastards’ll stop being so noisy.

Delaz

What do you have in mind? Weren’t you planning to court the Feds, using me and my fleet as bait? Are you going to drop the colony - or not?

Cima

Shut up, you old man!

Communications officer

The enemy’s breached the first line of defense. They’re closing in!

Cima

What?

Bait

Monsha, behind you!

Monsha

Hehe, I’m not that good!

Simmone

A direct hit on the left MS catapult.

Sinapus

Simmone, the time?

Simmone

The time? 4 minutes left.

Kou

Drop dead!

Kou

The colony - !

Gato’s fires an arm at the Dendrobium, finding its I-field engine and crushing it.

Kou

The I-field!

Gato

It’s too late.

Kou begins to rocket straight for the colony, avoiding Gato’s beam attacks and ignoring everything else.

Simmone

50 seconds left.

Cima

30 seconds to go.

Kou

It… passed the Critical Intercept Point…

Simmone

The colony’s headed. . . . . .for Earth.

Sinapus

It’s not over yet. We can’t let it fall on Jaburo.

Maurris

Captain. We’re receiving a message from the First Orbital Fleet.

Maurris

What!? They’re saying that the Delaz Fleet has ceased fire.

Sinapus

What do they mean by that?

Gato

Ceasefire!? This can’t be!

Federation ships positioned in front of the Critical Intercept Point assemble a huge array of mirrors.

Operator 1

This is G-21. We’ve finished laying out the upper passage of the Solar System.

Operator 2

This is left wing W-12 Unit. We’re having problems. We need reinforcements!

Staff

There are operational errors! Hurry! The enemy’s right in front of us!

Delaz

A Solar System…

Cima

That’s right. They had a card up their sleeve to stop the colony from falling. With just a little warming up… boom.

Delaz

Damn you! Do you still consider yourself a Jion?

Cima

This is how I survived! Despite being treated like shit by the brass who are living comfortably in Side 3! If there’s a good enough reason, I’ll switch sides!

The Neue Ziel approaches the Gwaden.

Gato

Your Excellency!

Delaz

What a pity. How foolish of me, trying to guide one without the soul of Jion.

Cima

Gee, I’ll go out of business the day I’m ’guided’ to some out-of-the-way place like Axis!

Gato

Cima, you ungrateful worm!

Cima

Don’t move, Gato. Accept your fate gracefully like a defeated general.

Gato

What?!

Cima

I don’t want to damage my gift to the Federation.

Delaz

Go on, Gato.

Delaz

Gato… carry out our cause. The colony is still in our hands.

Cima

What the -? Have you gone mad?

Gato

Step over my dead body and carry on.

Cima

Shut up!

Gato

Do you want to disgrace me with the worst punishment of space?

Cima

You fool! The Solar System is aiming at us! Y - You got to be kidding!

Gato

Your Excellency…

Delaz

Sieg Jion.

With a handgun, Cima executes Delaz. Gato explodes, firing an arm into the Gwaden’s bridge!

Ending song

EVERGREEN

Preview

The conspiracy proceeds, covering the sky of betrayal and a new age emerges. Will the colony finally reach the earth? A reunion of love and hate occurs amidst the battle. Now, what does Uraki see at the end of the war? An era fades away. UC 0083 is coming to an end.

Next: The Storm That Runs Through

END